OpenStudy (sasogeek):

A hemispherical tank of diameter 10m is filled by water issuin from a pipe of radius 20cm at 2m/s. calculate the time in minutes the time it takes to fill the tank

OpenStudy (anonymous):

Find the volume and then ...

OpenStudy (sasogeek):

i'm lost on this one...

OpenStudy (anonymous):

Um what is the volume of a hemisphere?

OpenStudy (bahrom7893):

half the volume of a sphere, come on saso, think...

OpenStudy (sasogeek):

2/3(pi r) cubed ?

OpenStudy (bahrom7893):

yup

OpenStudy (anonymous):

Good work, now what is the volume of a right cylinder?

OpenStudy (sasogeek):

\[\large \pi r^3 \]

OpenStudy (anonymous):

Don't you think there should be a variable for length too?

OpenStudy (sasogeek):

\[\large \pi r^3h \ \huge \text{ :D}\]

OpenStudy (anonymous):

Nopes ;) Hint: Volume = Area of the base \(\times\) Height

OpenStudy (sasogeek):

\[\large \text{facepalm :( } \huge \pi r^2h \]

OpenStudy (anonymous):

That's right saso :)

OpenStudy (sasogeek):

well what about the main question... i'm supposed to find the time it takes to fill the tank... i have volume, i have rate... what else?

OpenStudy (anonymous):

you are ready to put together the final pieces.

OpenStudy (sasogeek):

and what am i supposed to do with the radius of the pipe and i think the rate is missing something... isn't it supposed to be \(\large 2m^3s^-1 \)

OpenStudy (sasogeek):

\[\large 2m^3s^{-1} \]

OpenStudy (anonymous):

Here is the basic idea: Find the volume of the tank Find the volume of the water flowing through the pipe per second. Divide and find the time required in mints. Watch for units. I hope this should help.

OpenStudy (sasogeek):

thanks

OpenStudy (anonymous):

Glad to help saso :)